# Data-Efficient Operator Learning via Unsupervised Pretraining and In-Context Learning

### v1

@@ Abstract
Recent years have witnessed the promise of coupling machine learning methods and physical domain-specific insight for solving scientific problems based on partial differential equations (PDEs). However, being data-intensive, these methods still require a large amount of PDE data. This reintroduces the need for expensive numerical PDE solutions, partially undermining the original goal of avoiding these expensive simulations. In this work, seeking data efficiency, we design unsupervised pretraining and in-context learning methods for PDE operator learning. To reduce the need for training data with simulated solutions, we pretrain neural operators on unlabeled PDE data using reconstruction-based proxy tasks. To improve out-of-distribution performance, we further assist neural operators in flexibly leveraging in-context learning methods, without incurring extra training costs or designs. Extensive empirical evaluations on a diverse set of PDEs demonstrate that our method is highly data-efficient, more generalizable, and even outperforms conventional vision-pretrained models.
@@ 1 Introduction
Recent advancements in machine learning methodology have shown promise in solving partial differential equations (PDEs) (Han et al., 2018; Raissi et al., 2019; Kovachki et al., 2023; Li et al., 2021a, 2020; Lu et al., 2019; Krishnapriyan et al., 2023; Hansen et al., 2024). A significant development in this area is the concept of operator learning for PDEs. This approach differs from traditional neural network methods, which are restricted to fixed-dimension input and output, since neural operators focus on learning mappings between function spaces (Kovachki et al., 2023; Li et al., 2021a, 2020). As with other neural network methods, neural operators are recognized to be universal approximators for any continuous operator (Li et al., 2021b; Kovachki et al., 2023), enabling them to approximate any physical operator, including solution operators for various parametric PDE families. A solution operator is defined as a function that maps physical inputs (e.g., initial conditions) to output solutions. Previous work has shown that in simple settings neural operators can effectively capture complex, multi-scale dynamic processes (Li et al., 2021a, b; Takamoto et al., 2022; Subramanian et al., 2023).
@@
However, neural operator methods tend to suffer from a problem common to other DNN methods, namely the need for enormous quantities of data. The relatively limited availability of data is a common problem in science and engineering. High-fidelity numerical simulations are computationally costly or even infeasible for many applications (Sirignano et al., 2020). For example, an extreme-scale simulation of magnitude 7.0 earthquake at frequencies up to 10 Hz in the San Francisco requires 3600 Summit GPU nodes and 42.7 hour (McCallen et al., 2021).
Motivated by these data-efficiency challenges, recent work in machine learning, in particular natural language processing (NLP) and computer vision (CV), has focused on two general strategies for reducing the cost of collecting or generating labeled data for training: unsupervised (self-supervised) pretraining [^1] In this paper, we use the terms “unsupervised pretraining” and “self-supervised pretraining” interchangeably.; and in-context (few-shot) learning.
Unsupervised Pretraining.
@@
Unsupervised pretraining is widely used in CV and NLP to “pretrain” models. Such pretrained models have been shown to be highly data-efficient in downstream fine-tuning (Chen et al., 2020), and they can even become few-shot learners without any downstream data (Brown et al., 2020). In CV, people collect large amounts of natural images without any manual labels, and they then pretrain visual encoders with proxy tasks, such as Noise-Contrastive Estimation (NCE) (Oord et al., 2018), masked reconstruction (He et al., 2022), or rotation or jigsaw prediction (Noroozi and Favaro, 2016; Gidaris et al., 2018). In NLP, people typically pretrain models via next-word prediction or masked tokens (Brown et al., 2020; Devlin et al., 2018). However, unsupervised pretraining is still largely under-explored in Scientific Machine Learning (SciML). Therefore, our first question is:
@@
Q1: How can we design unsupervised pretraining for operator learning to reduce the data simulation costs?
In-Context Learning (ICL).
In-context learning (ICL) is a process by which a language model can make predictions on a test sample based on only a few demonstration examples provided in the input context (Brown et al., 2020; Logan IV et al., 2021). It is one of the most fascinating capabilities of large language models (LLMs), as this feature makes LLMs particularly versatile and adaptive to different downstream tasks—in particular for out-of-distribution (OOD) tasks, where only a limited number of downstream examples are needed, and without any fine-tuning cost. Recently, ICL has been introduced to operator learning. For example, during training, Yang et al. (2023a, b) enable the model to leverage a few provided support examples (or “demonstrations”) via self-attention, thus facilitating ICL when making predictions on OOD samples. However, this particular implementation of ICL, which introduces extra training costs as well as an explicit number of few-shot examples, deviates from the successful ICL practice in NLP, where language models are trained with simple token-prediction tasks and can easily scale up to any number of few-shot examples during inference. Therefore, our second question is:
Q2: How can we design a simple ICL for neural operators to achieve data-efficient OOD generalization?
@@
In this work, to achieve data efficiency in SciML, we propose unsupervised pretraining and ICL for neural operator learning. We overview our framework in Figure 1. To start, we propose two reconstruction-based proxy tasks. We demonstrate that with unsupervised pretraining, our neural operators not only improve counterparts trained with more simulated data, but also outperform off-the-shelf pretrained checkpoints from other popular domains (such as CV). Then, to further improve the data efficiency during OOD inference, we design a simple ICL strategy that is flexible and can scale up to any number of unseen supporting examples (“demos”). This approach also introduces zero overhead during training: one just keeps the standard training pipeline, and our ICL can be seamlessly plugged in (without further fine-tuning) for data-efficient OOD inference. In more detail, we summarize our main contributions:
@@
    We introduce unsupervised pretraining for data-efficient neural operator learning. We show that unsupervised pretraining on unlabeled PDE data can achieve better performance than models trained with more simulated data and publicly-available checkpoints pretrained on other benchmarks (such as CV), demonstrating the importance of pretraining on PDE domain data.
@@
    We propose a flexible ICL approach that can improve the OOD generalization of neural operators, without introducing any extra costs during training.
@@
    We provide a detailed empirical evaluation on diverse PDE benchmarks, demonstrating that we can achieve significant savings in PDE simulations and improve the performance of neural operators in OOD contexts.
@@ 2.1 Machine Learning for Scientific Modeling
There has been a long history of using machine learning-based methods to model (often nonlinear) physical scientific phenomena  (Lagaris et al., 1998, 2000; Chen and Chen, 1995a, b). A representative line of work is so-called physics-informed neural networks (PINNs) (Raissi et al., 2019; Zhu et al., 2019; Geneva and Zabaras, 2020; Gao et al., 2021; Ren et al., 2022), which try to incorporate physics in neural networks by including the differential form of the PDE as an additional physics loss regularization term. However, this paradigm is confined to specific PDE scenarios (e.g., fixed PDE coefficients) instead of being more physics-agnostic. Moreover, recent work has highlighted fundamental issues with PINN-based methods (Krishnapriyan et al., 2021; Edwards, 2022). On the other hand, operator learning methods, including Fourier Neural Operators (Li et al., 2021a, 2020; Kovachki et al., 2023) and Deep Operator Network (Lu et al., 2019), have achieved progress in approximating the solution operators of PDEs. Although these data-driven approaches show promise in learning PDE solutions, they (like many neural network-based methods) rely on vast quantities of high-fidelity labeled data. For operator learning, such data are usually computationally expensive to simulate (Raissi et al., 2019; Brandstetter et al., 2022; Zhang et al., 2023).
@@ 2.2 Unsupervised Pretraining and Foundation Models
Unsupervised (or self-supervised) pretraining is a key method in CV and NLP to achieve meaningful representations (Chen et al., 2020), data-efficient fine-tuning (Houlsby et al., 2019), and foundation models (Bommasani et al., 2021). In CV, contrastive learning learns meaningful features by distinguishing between similar (positive) and different (negative) samples (Oord et al., 2018; Tian et al., 2020; Chen et al., 2020; Misra and Maaten, 2020; He et al., 2020). Masked Autoencoder (MAE) (He et al., 2022) uses a reconstructive approach where parts of the input are masked and the model learns to predict masked parts. In NLP, among the most prominent works are large language models (LLMs) such as GPT (Brown et al., 2020; Radford et al., 2018, 2019) and BERT (Devlin et al., 2018), which leverage next token prediction for pretraining. The remarkable abilities (Wei et al., 2022) exhibited by foundation LLMs underscore the critical role of scale in achieving advanced capabilities that are not evident at smaller scales.
Similar directions also show progress in SciML.  Brandstetter et al. (2022) and Mialon et al. (2023) propose to create augmented views in the solution space via Lie Symmetries. Most notably, Subramanian et al. (2023) study the scaling behavior of supervised pretraining and OOD generalization, charting directions for foundational models for SciML.  Lanusse et al. (2023) targets learning astronomical foundation models with cross-modal contrastive learning.  McCabe et al. (2023) builds large task-agnostic models with a broad understanding of common physical behavior to serve as foundation models for SciML. [^2] As with others efforts motivated by the goal of foundational models for SciML, the data and scaling used in this paper certainly lack sufficient diversity to categorize the resulting models as “foundational.” However, our work marks the first extensive implementation of successful multiple nonlinear physics pretraining for spatiotemporal systems.
@@ 2.3 In-Context Learning
In-context learning is a promising paradigm in deep learning that helps deep networks generalize to unseen domains with a few in-context examples. Early works in computer vision seek to learn feature-level correspondence between the target and a few “shots,” such that models can generalize to open-set unseen objects (Fei-Fei et al., 2006; Zhang et al., 2020). In NLP, people find large language models (LLMs) are naturally few-shot learners (Brown et al., 2020), and thus tuning or optimizing prompts becomes extremely important to improve the in-context learning performance of LLMs (Zhou et al., 2022; Sordoni et al., 2023). More recently, within SciML, a different operator learning strategy, termed “in-context operator learning,” has been proposed (Yang et al., 2023a, b). During both training and inference, the neural operator is asked to make predictions by explicitly leveraging so-called “demo” examples (pairs of physical parameters and simulated solutions). This approach provides a balance between model generalization and addressing data scarcity in the scenario of OOD testing.
@@ 3.1 Unsupervised Pretraining
The core idea in unsupervised (or self-supervised) pretraining is to train a neural network with properly designed proxy tasks. These proxy tasks are unsupervised, i.e., they do not require labeled data, but they are designed to be highly related to the supervised learning objectives of interest. While popular in CV and NLP, unsupervised pretraining on unlabeled data has never been explored in PDE operator learning in SciML, mainly due to two unresolved questions: 1) What kinds of unlabeled data can we use to train neural operators? 2) How can we design proxy tasks for PDE data? We address each of these questions in turn.
3.1.1 Unlabeled Data
In SciML, neural operators are trained to map physical parameters and initial conditions to PDE solutions (Li et al., 2021a; Lu et al., 2019). Therefore, as a general approach to pretraining neural operators without labels (simulations), we will only use information in unlabeled PDE data.
Unlabeled PDE Data.
Our unlabeled PDE data is a broader concept of related inputs in modeling PDE systems. Let us consider the second-order linear differential equation as a general example. It is formulated as
@@ 3.1.1 Unlabeled Data
where $x\in\mathbb{R}^{n}$ represents physical space that varies in different systems ($n=3$ for 2D time-dependent PDEs); the coefficients (or physical parameters) $a_{ij},b_{i},c$ are known from the physical process; $u$ is the target solution; and $f$ denotes an external forcing function Nandakumaran and Datti (2020). We can consider two situations:
  • Time-independent equations: following previous works (Li et al., 2021a; Subramanian et al., 2023), our unlabeled PDE data include physical parameters ($a_{ij},b_{i},c$), forcing functions ($f$), and coordinates (grids of the discrete physical space).
  • Time-dependent equations (e.g., forecasting problems Takamoto et al. (2022); McCabe et al. (2023)): without simulating the temporal dynamics, our unlabeled PDE data include initial snapshot $u_{0}(x)$, which is sufficient for defining PDE systems.
There are two main reasons for pretraining only on unlabeled PDE data, as discussed below.
@@
Only generating unlabeled PDE data and snapshots without temporal dynamics will be much cheaper than simulating solutions, making our unsupervised pretraining highly feasible in practice. This pretraining strategy will be very data-efficient and can avoid the high computational cost of simulating complex time-dependent equations for massive high-fidelity labeled solutions (Table 8).
@@
Beyond the cheap generation, pretraining on unlabeled PDE data has the following benefits. First, regularization against overfitting in the low-data regime. When training on extremely low volumes of data, neural operators tend to overfit with poor generalization. Instead, unsupervised pretraining can strongly regularize the model towards better generalization, without the need for any early-stopping heuristics. Second, more meaningful representations. Pretraining on unlabeled PDE data can help models extract useful representations for subsequent operator learning. We defer our results and experimental details in Figure 4 in Sec. 4.1.
@@ 3.1.2 Proxy Tasks
To illustrate our general approach of constructing proxy tasks, we choose two variants of reconstruction as our core proxy tasks. In particular, we will input unlabeled PDE data to our neural operators, and after a decoder network we will force the output to be close to the input. We consider two perturbation variants (or augmented views).
@@
Masked autoencoders (MAEs) have been shown to be scalable self-supervised learners (He et al., 2022). The method is conceptually simple: remove a portion of the input data, and learn to predict the removed content. These methods enable training in NLP and CV of generalizable models containing over one hundred billion parameters (Devlin et al., 2018; Brown et al., 2020). Here, we investigate the potential of MAE for scientific modeling.
The motivation for using MAEs is that PDE dynamics are invariant to sparse sensing of the full field scientific data (Brunton et al., 2020; Erichson et al., 2020; Brunton et al., 2020). We enforce our model to learn sensor invariance via random masking, and we extract the invariant features over distorted views of the same unlabeled PDE data. The invariance to sparse sensing of scientific data will facilitate the robustness of the representations from MAE.
@@
Figure 2: Overview of unsupervised pretraining via MAE and super-resolution. During pre-training, in the input unlabeled PDE data, a random subset of spatial locations of (e.g., 70%) is masked out, followed by a Gaussian blur. After the encoder and decoder, the full set of input is required to be reconstructed.
@@
Super-resolution (SR) techniques have emerged as powerful tools for enhancing data resolution, improving the overall quality and fidelity of data representation, and retrieving fine-scale structures. SR is a task that involves recovering fine-scale data from corresponding coarse-grained data. It is also a popular task on PDE learning (Chung et al., 2023; Ren et al., 2023).
The motivation for using SR is that numerical solutions of PDEs are expected to exhibit invariance to filtering blur or different resolutions of initial conditions (Kochkov et al., 2021; Vinuesa and Brunton, 2022). For instance, in turbulence simulations, the traditional numerical methods always fail to model the expected physical phenomenon with low-resolution meshes due to substantial numerical errors. SR has emerged as a powerful tool for subgrid modeling of PDE dynamics, especially helping to capture the critical patterns of turbulence Maulik et al. (2019); Kochkov et al. (2021).
Therefore, we introduce SR as another proxy task. Specifically, we enforce the model to learn resolution-invariant features and to generate invariant features of the unlabeled PDE data which are immune to resolution and blur. To do so, we apply a Gaussian filter to blur the unlabeled PDE data, and the autoencoder will reconstruct the high-resolution unlabeled PDE data. Instead of applying a fixed blurring, we randomly sample the variance of the Gaussian filter from a certain range as augmentations.

### v2

@@ Abstract
Recent years have witnessed the promise of coupling machine learning methods and physical domain-specific insights for solving scientific problems based on partial differential equations (PDEs). However, being data-intensive, these methods still require a large amount of PDE data. This reintroduces the need for expensive numerical PDE solutions, partially undermining the original goal of avoiding these expensive simulations. In this work, seeking data efficiency, we design unsupervised pretraining for PDE operator learning. To reduce the need for training data with heavy simulation costs, we mine unlabeled PDE data without simulated solutions, and pretrain neural operators with physics-inspired reconstruction-based proxy tasks. To improve out-of-distribution performance, we further assist neural operators in flexibly leveraging in-context learning methods, without incurring extra training costs or designs. Extensive empirical evaluations on a diverse set of PDEs demonstrate that our method is highly data-efficient, more generalizable, and even outperforms conventional vision-pretrained models.
@@ 1 Introduction
Recent advancements in machine learning methodology have shown promise in solving partial differential equations (PDEs) [1, 2, 3, 4, 5, 6, 7, 8]. A significant development in this area is the concept of operator learning for PDEs. This approach differs from traditional neural network methods, which are restricted to fixed-dimension input and output, since neural operators focus on learning mappings between function spaces [3, 4, 5]. Like other neural network methods, neural operators are recognized to be universal approximators for any continuous operator [9, 3], enabling them to approximate any physical operator, including solution operators for various parametric PDE families. A solution operator is defined as a function that maps physical inputs to output solutions. Previous work has shown that in simple settings, neural operators can effectively capture complex, multi-scale dynamic processes [4, 9, 10, 11].
@@
However, neural operators tend to suffer from a problem common to other DNN methods, namely the need for enormous quantities of data. The limited availability of data is common in science and engineering. High-fidelity numerical simulations are computationally costly or even infeasible for many applications [12]. For example, an extreme-scale simulation of magnitude 7.0 earthquake at frequencies up to 10 Hz in San Francisco requires 3600 Summit GPU nodes and 42.7 hour [13].
@@
Motivated by this data-efficiency challenge, recent works, particularly in natural language processing (NLP) and computer vision (CV), has focused on unsupervised (or self-supervised) pretraining [^1] In this paper, we use the terms “unsupervised pretraining” and “self-supervised pretraining” interchangeably. to reduce the cost of collecting or generating labeled data. Such pretrained models have been shown to be highly data-efficient in downstream fine-tuning [14, 15, 16], and they can even become few-shot learners without any downstream data [17]. In CV, researchers collect large amounts of natural images without any manual labels, and then pretrain visual encoders with proxy tasks, such as Noise-Contrastive Estimation (NCE) [18], masked reconstruction [19], rotation and jigsaw prediction [20, 21]. In NLP, people typically pretrain models via next-word prediction or masked tokens [17, 22].
However, unsupervised pretraining is still largely underexplored in Scientific Machine Learning (SciML). Therefore, our core question is:
@@
How to design unsupervised pretraining for operator learning to reduce the data simulation costs?
@@
In this work, we resort to unsupervised pretraining for neural operator learning to achieve data efficiency in SciML. We overview our framework in Figure 1. To start, we define unlabeled data for PDEs, which avoids heavy computation costs for simulating PDE solutions. We propose two physics-inspired reconstruction-based proxy tasks and pretrain neural operators on unlabeled PDE data. We demonstrate that with unsupervised pretraining, our neural operators not only improve counterparts trained with more simulated data, but also outperform off-the-shelf pretrained checkpoints from other popular domains (such as CV) that are ready for fine-tuning. Then, to further improve the data efficiency during out-of-distribution (OOD) inference, we design a simple in-context learning (ICL) strategy [17, 23, 24, 25, 26]. This approach introduces zero overhead during training: one just maintains the standard training pipeline, and it can be seamlessly plugged in for OOD inference without further fine-tuning. In more detail, we summarize our main contributions:
@@
    We introduce unlabeled PDE data and unsupervised pretraining for data-efficient neural operator learning. We show that our method can achieve better performance than models trained with more simulated PDE solutions, or fine-tuned from public checkpoints pretrained on other benchmarks, demonstrating the importance of unsupervised pretraining on domain-specific PDE data.
@@
    We propose an ICL approach to improve the OOD generalization of neural operators, which is flexible and can scale up to a large number of unseen supporting examples (“demos”).
@@
    We provide detailed empirical evaluations on both diverse PDE benchmarks and also real-world scenarios, demonstrating that we can achieve both strong forward modeling performance and significant savings in PDE simulations.
@@ 2.1 Machine Learning for Scientific Modeling
There has been a long history of using learning-based methods to model physical scientific phenomena  [27, 28, 29, 30]. A representative line of work is so-called physics-informed neural networks (PINNs) [2, 31, 32, 33, 34], which try to incorporate physics in neural networks by including the differential form of the PDE as an additional physics loss regularization term. However, this paradigm is confined to specific PDE scenarios (e.g., fixed PDE coefficients) instead of being more physics-agnostic. Moreover, recent work has highlighted fundamental issues with PINN-based methods [35, 36]. On the other hand, operator learning methods, including Fourier Neural Operators [4, 5, 3] and Deep Operator Network [6], have achieved progress in approximating the solution operators of PDEs. Although these data-driven approaches show promise in learning PDE solutions, they (like many neural network-based methods) rely on vast quantities of high-fidelity labeled data. For operator learning, such data are usually computationally expensive to simulate [2, 37, 38].
@@ 2.2 Unsupervised Pretraining and Foundation Models
Unsupervised (or self-supervised) pretraining is a key method in CV and NLP to achieve meaningful representations [14], data-efficient fine-tuning [39], and foundation models [40]. In CV, contrastive learning learns meaningful features by distinguishing between similar (positive) and different (negative) samples [18, 41, 14, 42, 15]. Masked Autoencoder (MAE) [19] uses a reconstructive approach where parts of the input are masked and the model learns to predict masked parts. In NLP, among the most prominent works are large language models (LLMs) such as GPT [17, 43, 44] and BERT [22], which leverage next token prediction for pretraining. Similar directions also show progress in SciML.  [37] and [45] propose to create augmented views in the solution space via Lie Symmetries. Most notably, [11] study the scaling behavior of supervised pretraining and OOD generalization, charting directions for foundational models for SciML.  [46] targets learning astronomical foundation models with cross-modal contrastive learning.  [47] builds large task-agnostic models with a broad understanding of common physical behavior to serve as foundation models for SciML.
@@ 2.3 In-Context Learning
In-context learning is a promising paradigm that helps deep networks generalize to unseen domains with a few in-context examples. Early works in CV seek to learn feature-level correspondence between the target and a few “shots,” such that models can generalize to open-set unseen objects [48, 49]. In NLP, people find large language models (LLMs) are naturally few-shot learners [17], and thus tuning or optimizing prompts becomes extremely important to improve the in-context learning performance of LLMs [50, 51]. More recently, within SciML, a different operator learning strategy, termed “in-context operator learning,” has been proposed [24, 25, 26]. During both training and inference, the neural operator is asked to make predictions by explicitly leveraging a predefined number of so-called “demo” examples (pairs of physical parameters and simulated solutions). This approach provides a balance between model generalization and addressing data scarcity in the scenario of OOD testing.
@@ 3.1 Unsupervised Pretraining
The core idea in unsupervised (or self-supervised) pretraining is to train a neural network with properly designed proxy tasks. These proxy tasks do not require labeled data, but they are designed to be highly related to the supervised learning objectives of interest. While popular in CV and NLP, unsupervised pretraining on unlabeled data has never been explored in PDE operator learning in SciML, mainly due to two unresolved questions: 1) What kinds of unlabeled data can we use to train neural operators? 2) How to design proxy tasks for PDE data? We address these questions below.
3.1.1 Unlabeled PDE Data
How to Define Unlabeled PDE Data?
In general, when a neural operator is trained on PDE datasets [4, 10, 11], it learns to map inputs (physical parameters, coordinates, forcing functions, initial conditions, etc.) to PDE solutions. Therefore, given a set of PDE data (collected via simulations or observations), its unlabeled version is defined as the one without PDE solutions. Our unlabeled PDE data is a broader concept of related inputs in modeling PDE systems. Let us consider the second-order linear differential equation as a general example. It is formulated as
@@ 3.1.1 Unlabeled PDE Data
where $x\in\mathbb{R}^{n}$ represents physical space that varies in different systems (e.g. $n=3$ for 2D time-dependent PDEs); the coefficients (or physical parameters) $a_{ij},b_{i},c$ are known from the physical process; $u$ is the target solution; and $f$ denotes an external forcing function [52]. We can consider two situations where solutions are unavailable:
  • Time-independent equations: following [4, 11], our unlabeled PDE data include physical parameters ($a_{ij},b_{i},c$), forcing functions ($f$), and coordinates (grids of the discrete physical space).
  • Time-dependent equations (e.g., forecasting problems [10, 47]): without simulating the temporal dynamics, our unlabeled PDE data only include initial snapshot $u_{0}(x)$, which is sufficient for defining PDE systems.
For concrete examples of different PDEs and unlabeled data that we will study, please see Appendix A. There are two main reasons for pretraining only on unlabeled PDE data, as discussed below.
@@
One critical reason that leads to expensive computational costs when collecting PDE data is the time marching scheme [53, 54] in numerical simulation. However, only generating unlabeled PDE data and snapshots without temporal dynamics will be much cheaper than simulating solutions (Table 3), making our unsupervised pretraining highly feasible in practice. Our pretraining strategy will be very data-efficient, and can avoid the heavy computational cost of simulating complex time-dependent equations for massive high-fidelity labeled solutions.
@@
Beyond the cheap generation, pretraining on unlabeled PDE data has the following benefits. First, regularization against overfitting. Unsupervised pretraining can strongly regularize the model towards better generalization. Second, faster convergence. Pretraining on unlabeled PDE data will provide neural operators with domain-adapted initializations and can accelerate training speed. Third, meaningful representations. Pretraining on unlabeled PDE data can help models extract useful representations for subsequent operator learning. We defer our results and experimental details in Figure 4 in Sec. 4.1.
@@ 3.1.2 Proxy Tasks
To illustrate our general approach of constructing proxy tasks, we choose two variants of reconstruction as our core proxy tasks. In particular, we will input unlabeled PDE data to our neural operators, and after a decoder network, we will force the output to be close to the input. We consider two perturbation variants (or augmented views). They are inspired by real-world settings when people collect scientific data, and they are important invariances we need to introduce to SciML models.
@@
Masked autoencoders (MAEs) have been shown to be scalable self-supervised learners [19]. The method is conceptually simple: remove a portion of the input data, and learn to predict the removed content. These methods enable training in NLP and CV of generalizable models containing over one hundred billion parameters [22, 17]. Here, we investigate the potential of MAE for scientific modeling.
Motivation: PDE dynamics are invariant to sparse sensing of the full field scientific data It is very common [55, 56, 57] that scientific data to be collected from sparse sensors, and people need to reconstruct or generate data for domains without sensors. We enforce our model to learn sensor invariance via random masking, and we extract the invariant features over distorted views of the same unlabeled PDE data. The invariance to sparse sensing of scientific data will facilitate the robustness of the representations from MAE.
@@
Figure 2: Overview: unsupervised pretraining via MAE and super-resolution. During pre-training, in the input unlabeled PDE data, a random subset (e.g., 70%) of spatial locations are masked, followed by a Gaussian blur. After the encoder and decoder, the full set of input is required to be reconstructed.
@@
Super-resolution (SR) techniques have emerged as powerful tools for enhancing data resolution, improving the overall quality and fidelity of data representation, and retrieving fine-scale structures. SR is a task that involves recovering fine-scale data from corresponding coarse-grained data. It is also a popular task on PDE learning [58, 59].
Motivation: Numerical solutions of PDEs are expected to exhibit invariance to filtering blur or different resolutions of inputs [60, 61]. For instance, in turbulence simulations, the traditional numerical methods always fail to model the expected physical phenomenon with low-resolution meshes due to substantial numerical errors. SR has emerged as a powerful tool for subgrid modeling of PDE dynamics, especially helping to capture the critical patterns of turbulence [62, 60].
Therefore, we introduce SR as another proxy task. Specifically, we enforce the model to learn resolution-invariant features and to generate invariant features of the unlabeled PDE data which are immune to resolution and blur. Blurry snapshots often occur when the resolution is too low to accurately represent the details in the original content. To do so, we apply a Gaussian filter to blur the unlabeled PDE data, and the autoencoder will reconstruct the high-resolution input with fine-grained details. Instead of applying a fixed blurring, we randomly sample the variance of the Gaussian filter from a certain range as augmentations.